OpenStudy (anonymous):

Calculate the earth's speed in its orbit around the sun,in m/s, km/h and mi/h.

OpenStudy (anonymous):

we should equate centrifugal force with gravitational force \[mv^2/r=GMm/r^2\] \[v=\sqrt{GM/r}\] where M= mass of sun m=mass of earth r distance b/w them